OpenStudy (anonymous):

3P(0)=P(0)e^15%*t i need to find the time(years) it takes for that amount of money to triple if compounded continuously at a rate of 15%

OpenStudy (anonymous):

P is the principal and the amt is 3P annual rate of interest is 15% let t be the no of yrs then we have 3P=P(1+.15)^t 3=(1+.15)^t or 3= 1.15^t taking log to the base 10 both sides we hav log 3= t * log 1.15 or t = log 3 / log 1.15

OpenStudy (anonymous):

t=0.4771/.06 =7.951 = appx 8 yrs

OpenStudy (kropot72):

The formula for continuous compounding is as follows: \[A=Pe ^{rt}\] where A is the amount after t years, P is the principal and r is the annual interest rate expressed as a decimal. e is the exponential number. Using the formula in your question, but expressing the interest rate as a decimal we get \[3P _{0}=P _{0}e ^{0.15t}\ ............(1)\] Simplifying equation (1) gives \[3=e ^{0.15t}\ .................(2)\] Taking natural logs of both sides of equation (2) we get \[\ln 3=0.15t\ ...............(3)\] and we find the time t from \[t=\frac{\ln 3}{0.15}=7.324\ years\]